OpenStudy (anonymous):

write an algebraic expression for each verbal expression. r=2,s=5,and t=1/2 1. t times the sum of four times s and r 2.r to the fifth power decreased by t

OpenStudy (anonymous):

t x 4(s+r) = 1/2 x 4(2+5) = 1/2 x 4(7) = 1/2 x 28 = 14 r^5 - t = 2^5 - 1/2 = 25 - 1/2 = 24.5

OpenStudy (anonymous):

I don't think the previous replies are correct. I'm getting something different. 1) t(4s+r) = 4st + rt = 4 x 5 x 1/2 + 2 x 1/2 = 10 + 1 = 11 2) r^5 - t = 2^5 - 1/2 = 32 - 1/2 = 31.5

OpenStudy (anonymous):

haha my bad. I definitely got some numbers mixed up. Thanks math_tutor!

OpenStudy (anonymous):

No problem! It's easy to do that with numbers. LOL